Good evening, I'm having trouble determining how to find the x-intercepts of the following function:
y = 2(x-3)^2 - 23
I found the y-intercept by setting x to zero:
y = 2(0-3)^2 - 23
y = 18-23
y = -5
However, I know I'm not finding the x-intercepts correctly, as I'm only coming up with one x-intercept using the quadratic equation:
x = [12 +/- sqrt(144 - 4(2)(18))] / 2(2)
x = (12 +/- 0)/4
x = 3
Could you let me know how I should be finding the x-intercept? Thanks so much for sharing your knowledge with us students :)

Answer by mananth(16949)

y = 2(x-3)^2 - 23
for x intercept y =0
2(x-3)^2=23
(x-3)^2=23/2
take the square root
x-3 = +/- sqrt(23/2)
x1= 3+sqrt(23/2)
6.39
x2=3-sqrt(23/2)
x2=-0.39
,,,
your intercepts are ready!
m.ananth@hotmail.ca